Air initially at 100 0F and 100 psia and occupying a volume of 0.5 ft3 undergoes a reversible non-flow constant temperature process such that the final pressure becomes 20 psia. find the work done, ft.lbf.
1
Expert's answer
2020-05-14T14:37:42-0400

"T_1= 100 ^0F , P_1= 100 psi,V_1=0.5 ft^3 , P_2= 20 ,"


at constant temperature

"W= P_1V_1ln(\frac{P_2}{P_1})= 100 \times 0.5 ln(\frac{100}{20})= 80.47" ftlbf